# Multi-Modal Framing Analysis of News

### v1

@@ Abstract
Arnav Arorathanks:   denotes equal contribution      Srishti Yadav [^1] https://www.foxnews.com/media/chicago-migrants-return-venezuela-winter-nothing-here Maria Antoniak      Serge Belongie      Isabelle Augenstein University of Copenhagen {aar, srya, maan, s.belongie, augenstein}@di.ku.dk
@@ 1 Introduction
News media often employ images alongside text to reinforce or contrast the intended frame, leveraging the affective and cognitive impact of visuals Cope et al. (2005); Keib et al. (2018); Grabe and Bucy (2009); Geise et al. (2025). An example can be seen in Figure 1.111https://www.foxnews.com/media/chicago-migrants-return-venezuela-winter-nothing-here The image depicts migrants in poor living conditions, talking about quality of life and depicting a lack of resources. The article text, on the other hand, also talks about the legal aspects of immigration and economics of their sustenance. Such differences in portrayal can affect readers’ perception of the crisis and their induced emotions. Further, while communication in text is more explicit through linguistic framing, images encode frames in more subtle and implicit ways, requiring sophisticated interpretation models to capture their meaning Aiello and Parry (2019). As such, when conducting media analysis, the communicated framing across both the image and the text should be considered.
@@ 2.2 Integrative Framing Analysis
Integrative framing analysis is when both images and text are observed separately, but the results are integrated to form a more complete picture of framing analysis (Dan, 2017). Although there is broad agreement that visual and verbal elements should be studied side by side (Coleman, 2010; Rose, 2022), relatively few studies have effectively combined the two. Researchers often focus on textual framing Matthes and Kohring (2008) or visual framing Rodriguez and Dimitrova (2011) as separate domains, even though media messages typically blend both. Geise et al. (2013) and Coleman (2010) emphasize that analyzing visuals in isolation can obscure how words shape viewers’ interpretations-and vice versa. As computational tools become increasingly sophisticated and as scholarly interest in visual framing continues to grow, integrative approaches are poised to become more common. By integrating these modalities, scholars can better observe whether text and visuals reinforce each other or present divergent frames.
@@ 4.2 Models and Prompts
For text annotations, we use Mistral-7B. For images, we use the multi-modal model Pixtral-12B. We use the vLLM library for optimized memory usage and high-throughput inference, allowing us to conduct analysis on the entire dataset in a few days, demonstrating our approach’s scalability. See Appendix D for full modelling details.
@@ 4.3 Model Evaluation
To assess the quality of output frames from our approach, and compare against prior work, we evaluate our text frame extraction model on the existing large scale benchmark Media Frames Corpus (Card et al., 2015). This corpus provides frame labels by each annotator for over 32,000 articles from US news on topics like immigration, smoking, and same-sex marriage. We take the union of all frame labels assigned by the annotators and assign the up to the top 3 most frequent frames assigned by the annotators as the label for an article. We run our text framing analysis model on this dataset, providing the article of the text while allowing the model to generate multiple frames per article. To assess model performance, we calculate the intersection between the sets of model-annotated frames and human-annotated frames for each article in the dataset. 89.36% of the articles had a non-zero intersection, with at least one overlapping frame label, demonstrating that the model outputs accurate frames in most cases. Our model received a micro averaged F1 score of 0.44, with an averaged precision of 0.4 and averaged recall of 0.5. We include the scores per frame label in Appendix F. For prediction across 15 labels and a task as subjective as framing, we believe the model performs quite well. We further did qualitative assessment of model outputs by going through the model explanations for the predicted frames and found the model to be competent.
@@
For images, there is no existing benchmark with frame labels. Two of the authors of this study manually annotated 600 images for generic frames across a stratified (time-period and topics) sample of the dataset. We modify generic frame annotation guidelines released by Mendelsohn et al. (2021), adapting them to the image annotation setting. We set up a multi-label classification platform (as seen in Figure 12) where annotators are allowed to select one or more frames from the list given an image. Framing analysis is a highly subjective task, as is well established by prior work in communication studies as well as NLP. For images, the subjectivity increases substantially, given the limited amount of context available and requiring cultural and conceptual familiarity more so than needed for textual framing analysis (Geise and Baden, 2015). We calculate the proportional overlap in frames between annotators, counting a non-zero intersection of frame annotations as 1 and zero intersections as 0. The proportional overlap between the annotators is 85%, demonstrating the high percentage of instances with overlapping frame labels. Using this dataset, we create an evaluation set for our vision model. We take the union of the frame predictions for the two annotators and calculate the intersection with the model predictions for a particular image. The proportion of non-zero intersection instances between the image and human annotations is 84.2%, i.e., at least one correctly predicted frame most of the time.
@@ Appendix A Source selection
Micro avg 0.40 0.50 0.44
Macro avg 0.35 0.52 0.40
Wt. avg 0.38 0.50 0.41
Samples avg 0.44 0.51 0.44
@@ Appendix G Additional Plots
prompt_entity = """
Your task is to identify the main subject or entity in the image that is central to the image. Entities are people or organizations that play a central role in the image. If there are multiple entities, choose the one that is central to the image. if there are no central or clearly identifiable entities, choose "None".
@@
For the image above, provide the name of the entity, sentiment towards the entity, and reasoning for the chosen sentiment.
@@
Policy prescription and evaluation - discussion of specific policies aimed at addressing problems. In the image, we can see things including but not limited to discussions on rule, rule making bodies, people in formal settings - such as boardrooms or legislative halls - actively debating, and reviewing policy drafts or proposals. You might see official charts, graphs, or official documents. People in formal attire with no other information should not be classified as policy prescription and evaluation.

### v2

@@ Abstract
Arnav Arorathanks:   denotes equal contribution      Srishti Yadav11footnotemark: 1 Maria Antoniak      Serge Belongie      Isabelle Augenstein University of Copenhagen {aar, srya, maan, s.belongie, augenstein}@di.ku.dk
@@ 1 Introduction
News media often employ images alongside text to reinforce or contrast the intended frame, leveraging the affective and cognitive impact of visuals Cope et al. (2005); Keib et al. (2018); Grabe and Bucy (2009); Geise et al. (2025). An example can be seen in Figure 1. The image depicts protesters with signs with police presence around, using the public opinion and security framing. The article text, on the other hand, talks about the quality of life of migrants, economic implications, crime, and policy framing. Such differences in portrayal can affect readers’ perception of the crisis and their induced emotions. Further, while communication in text is more explicit through linguistic framing, images encode frames in more subtle and implicit ways, requiring sophisticated interpretation models to capture their meaning Aiello and Parry (2019). As such, when conducting media analysis, the communicated framing across both the image and the text should be considered.
@@ 2.2 Integrative Framing Analysis
Integrative framing analysis is when both images and text are observed separately, but the results are integrated to form a more complete picture of framing analysis (Dan, 2017). Although there is broad agreement that visual and verbal elements should be studied side by side (Coleman, 2010; Rose, 2022), relatively few studies have effectively combined the two. Researchers often focus on textual framing Matthes and Kohring (2008) or visual framing Rodriguez and Dimitrova (2011) as separate domains, even though media messages typically blend both. Geise et al. (2013) and Coleman (2010) emphasize that analyzing visuals in isolation can obscure how words shape viewers’ interpretations—and vice versa. As computational tools become increasingly sophisticated and as scholarly interest in visual framing continues to grow, integrative approaches are poised to become more common. By integrating these modalities, scholars can better observe whether text and visuals reinforce each other or present divergent frames.
@@ 4.2 Models and Prompts
For text annotations, we use Mistral-7B. For images, we use Pixtral-12B (multi-modal model). We use the vLLM library for optimized memory usage and high-throughput inference, allowing us to conduct analysis on the entire dataset in a few days, demonstrating our approach’s scalability. See Appendix D for full modelling details.
@@ 4.3 Model Evaluation
To assess the quality of output frames from our approach, and compare against prior work, we evaluate our text frame extraction model on the existing large scale benchmark Media Frames Corpus (Card et al., 2015). This corpus provides frame labels by each annotator for over 32,000 articles from US news on topics like immigration, smoking, and same-sex marriage. We take the union of all frame labels assigned by the annotators and assign the up to the top 3 most frequent frames assigned by the annotators as the label for an article. We run our text framing analysis model on this dataset, providing the article of the text while allowing the model to generate multiple frames per article. To assess model performance, we calculate the intersection between the sets of model-annotated frames and human-annotated frames for each article in the dataset. 95.7% of the articles had a non-zero intersection, with at least one overlapping frame label, demonstrating that the model outputs accurate frames in most cases. Our model received a micro averaged F1 score of 0.5, with an averaged precision of 0.42 and averaged recall of 0.62. We include the scores per frame label in Appendix F. For prediction across 15 labels and a task as subjective as framing, we believe the model performs quite well. We further did qualitative assessment of model outputs by going through the model explanations for the predicted frames and found the model to be competent.
@@
For images, there is no existing benchmark with frame labels. Two of the authors of this study manually annotated 600 images for generic frames across a stratified (time-period and topics) sample of the dataset. We modify generic frame annotation guidelines released by Mendelsohn et al. (2021), adapting them to the image annotation setting. We set up a multi-label classification platform (as seen in Figure 12) where annotators are allowed to select one or more frames from the list given an image. Framing analysis is a highly subjective task, as is well established by prior work in communication studies as well as NLP. For images, the subjectivity increases substantially, given the limited amount of context available and requiring cultural and conceptual familiarity more so than needed for textual framing analysis (Geise and Baden, 2015). We then calculate the proportional overlap in frames between annotators, counting a non-zero intersection of frame annotations as 1 and zero intersections as 0. The proportional overlap between the annotators is 85%, demonstrating the high percentage of instances with overlapping frame labels. Using this dataset, we create an evaluation set for our vision model and get predictions from our pixtral model. Since ’Sports’ news are mostly labelled as ’cultural identity’, we remove them (8%of them) for our calculation. We take the union of the frame predictions for the two annotators and calculate the intersection with the model predictions for a particular image. The proportion of non-zero intersection instances between the image and human annotations is 84.2%, i.e., at least one correctly predicted frame most of the time.
@@ Appendix A Source selection
micro avg 0.42 0.62 0.50
macro avg 0.39 0.58 0.45
weighted avg 0.45 0.62 0.51
samples avg 0.46 0.63 0.51
@@ Appendix G Additional Plots
prompt_entity = """Your task is to identify the main subject or entity in the image that is central to the image. Entities are people or organizations that play a central role in the image. If there are multiple entities, choose the one that is central to the image. if there are no central or clearly identifiable entities, choose "None".
@@
For the image above, provide the name of the entity, sentiment towards the entity, and reasoning for the chosen sentiment."""
@@
Policy prescription and evaluation - discussion of specific policies aimed at addressing problems. In the image, we can see things including but not limited to discussions on rule, rule making bodies, people in formal settings such as boardrooms or legislative halls - actively debating, and reviewing policy drafts or proposals. You might see official charts, graphs, or official documents. People in formal attire with no other information should not be classified as policy prescription and evaluation.